OpenStudy (anonymous):

integrate (sinhx + 2 coshx)/(sinhx +coshx)

OpenStudy (anonymous):

should i let u= sinhx +coshx ?

OpenStudy (irishboy123):

sinh + cosh = ??

OpenStudy (anonymous):

1??

OpenStudy (loser66):

multiple both numerator and denominator by (cosh (x) -sinh (x))

jimthompson5910 (jim_thompson5910):

Hints \[\Large \sinh(x) = \frac{e^x-e^{-x}}{2}\] \[\Large \cosh(x) = \frac{e^x+e^{-x}}{2}\] http://math2.org/math/trig/hyperbolics.htm

OpenStudy (loser66):

Then denominator = \(cosh^2 (x) - sinh^2(x) =1\) you have just numerator, right? What do you have?

OpenStudy (anonymous):

wait a moment please

OpenStudy (anonymous):

@Loser66 the numerator = \[2\cosh^2x -\sinh^2x -coshxsinhx\]

OpenStudy (loser66):

yup, and 2 cosh^2 = cosh^2 + cosh^2 , right? again, cosh^2 - sinh^2 =1, Now, what do you get?

OpenStudy (anonymous):

(sinh^2 + 1)(sinh^2+1)

jimthompson5910 (jim_thompson5910):

Use these identities \[\Large {\color{red}{\sinh(x)}} = {\color{red}{\frac{e^x-e^{-x}}{2}}}\] \[\Large {\color{blue}{\cosh(x)}} = {\color{blue}{\frac{e^x+e^{-x}}{2}}}\] To make these substitutions \[\Large \frac{\sinh(x)+2\cosh(x)}{\sinh(x)+\cosh(x)}\] \[\Large \frac{{\color{red}{\sinh(x)}}+2*{\color{blue}{\cosh(x)}}}{{\color{red}{\sinh(x)}}+{\color{blue}{\cosh(x)}}}\] \[\Large \frac{{\color{red}{\frac{e^x-e^{-x}}{2}}}+2*{\color{blue}{\frac{e^x+e^{-x}}{2}}}}{{\color{red}{\frac{e^x-e^{-x}}{2}}}+{\color{blue}{\frac{e^x+e^{-x}}{2}}}}\] Now simplify. After simplifying, the integration should be much easier.

OpenStudy (loser66):

If you want to follow Mr. JIm's way, it's ok. There are many ways to get the answer. I finish my way :) \(2 cosh^2 (x) -sinh^2 (x) -cosh(x) sinh(x) \\= cosh^2(x) +cosh^2(x) -sinh^2(x) -cosh(x)sinh(x)\\=cosh^2(x) -1 -cosh(x) sinh(x)\) Now take integral term by term. 1 and cosh(x) sinh(x) are easy, right? Just the first element. We have \(cosh(2x) = 2cosh^2(x) -1\), then \(cosh^2 (x) = cosh(2x)/2+1/2\) Every thing is done. right? Of course, you have to put them in neat. Ok?

OpenStudy (anonymous):

why 1 is negative, arent cosh^2 x=1+sinh^2 x

OpenStudy (loser66):

yes, my bad :)

OpenStudy (solomonzelman):

you have to be careful there, when applying the same rules of reg trig functions to hyperbolic...

OpenStudy (solomonzelman):

\(2\cosh^2(x)-1=\cosh(2x)\) is not an identityy...

OpenStudy (loser66):

@SolomonZelman Yes, I read it before using it https://en.wikipedia.org/wiki/Hyperbolic_function

OpenStudy (solomonzelman):

omg, thatlink is so nice... tnx 66 !

OpenStudy (solomonzelman):

yeah, it is an identity, my badd

OpenStudy (solomonzelman):

purposefully

OpenStudy (solomonzelman):

well, if you want, Loser66 :) ... why 66, btw?

OpenStudy (solomonzelman):

Age changes, username on os doesn't ... anyway ... that is reasonable however...

OpenStudy (solomonzelman):

back to problem, won't distract...

OpenStudy (solomonzelman):

eunna, did you simplify as directed by Jim Thompson?

OpenStudy (anonymous):

i try to simplify, but im bad at simplifying, so the answer is weird.. :')

OpenStudy (solomonzelman):

\({\LARGE \displaystyle \frac{ \frac{e^x-e^{-x}}{2} + 2\times \frac{e^x+e^{-x}}{2} }{ \frac{e^x-e^{-x}}{2}+ \frac{e^x+e^{-x}}{2} }}\) \({\LARGE \displaystyle \frac{ \frac{e^x-e^{-x}}{2} + \frac{2e^x+2e^{-x}}{2} }{ \frac{e^x-e^{-x}}{2}+ \frac{e^x+e^{-x}}{2} }}\)

OpenStudy (solomonzelman):

\({\LARGE \displaystyle \frac{ \frac{e^x-e^{-x}+2e^x+2e^{-x}}{2} }{ \frac{e^x-e^{-x}+e^x+e^{-x}}{2} }}\)

OpenStudy (solomonzelman):

show me please, what do you get after simplifying this...

OpenStudy (solomonzelman):

you can multiply times two on top and bottom for simplicity \({\Large \displaystyle \frac{ e^x-e^{-x}+2e^x+2e^{-x} }{e^x-e^{-x}+e^x+e^{-x} }}\)

OpenStudy (solomonzelman):

please continue

OpenStudy (anonymous):

give me a moment

OpenStudy (solomonzelman):

k

OpenStudy (anonymous):

\[\frac{ e^x -e ^{-x}+2e^x+2e-x}{ \frac{ 2 }{ e^x } }\]

OpenStudy (solomonzelman):

when yo add everything on top and bottom, \({\Large \displaystyle \frac{ 3e^x+e^{-x}}{2e^x }}\)

OpenStudy (solomonzelman):

\({\Large \displaystyle \frac{ 3e^x}{2e^x }+\frac{ e^{-x}}{2e^x }}\) \({\Large \displaystyle \frac{ 3}{2 }+\frac{ 1}{2e^{2x} }}\)

OpenStudy (solomonzelman):

so really that is what you are integrating ... (shouldn't be so hard to do)

OpenStudy (anonymous):

i got different answer for both method

OpenStudy (anonymous):

thank you guys :)